OpenStudy (anonymous):

4. Use the equation you have written above determine the cost for each bracelet. Show the algebraic steps that it takes to find the answer. Provide your conclusion.

OpenStudy (anonymous):

Provide equation

OpenStudy (anonymous):

alright hold on i already solved the equation

OpenStudy (dan815):

ok

OpenStudy (anonymous):

3. The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Define your variable and write an equation that models the cost of each bracelet.

OpenStudy (anonymous):

this is what i solved now they want me to show the algebraic steps to the question i just posted that i solved I am not good at algebra

OpenStudy (anonymous):

54=8x + 6

OpenStudy (anonymous):

8x+6=54 8x-6=8x 8x/8=x 54-6=48 48/8=6 x=$6 per bracelet but i need to kow what that mean with this question. 4. Use the equation you have written above determine the cost for each bracelet. Show the algebraic steps that it takes to find the answer. Provide your conclusion.

OpenStudy (anonymous):

i got that but what i am trying to get is what they want me to do with this question i just posted. it supposely goes with that other question

OpenStudy (anonymous):

Nothing. It's just a simple break down. You have to provide steps like "Let X be the variable representing cost of a bracelet" "8 bracelets then cost 8X". "One time shipping cost is 6". Sum of all this is 54

OpenStudy (anonymous):

so that is what i write for the question 4?

OpenStudy (anonymous):

Yes, kind of like a line by line walkthrough

OpenStudy (anonymous):

So arsalanaijaz do i answer the question like this x represents the variable cost of the bracelets. 8 bracelets cost 8x One time shipping fee cost is $6. Sum of all this is $54 Is that right or wrong

OpenStudy (anonymous):

That should do it. Then show the working and Conclude that x = something, implies cost of bracelet is something

OpenStudy (anonymous):

what you mean show the work

OpenStudy (anonymous):

didnt i do the work on question 3 so i have to redo the work i did on question on question 4

OpenStudy (anonymous):

Oh, that's another question! Yup, you're right.

OpenStudy (anonymous):

so show the work i did on 3 on number 4

OpenStudy (anonymous):

You did math in 3. Do the english explanation on 4

OpenStudy (anonymous):

oh ok so i just have to write what i wrote to you and thats all no solving the problem again

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you might have other questions i need help with if i do would you be able to help me out

OpenStudy (anonymous):

yup, if i'm around

OpenStudy (anonymous):

oh ok i do have some question i dont understand

OpenStudy (anonymous):

alright arsalanaijaz you there cause i have on the next question i dont get how to solve it